Dhorkas owns a small restaurant. She has labourer who she pays GHC 12,000 per year, pays annual rent of GHC 5,000 for her shop, and spends GHC 20,000 per year on the materials she uses for the preparation of her meals. She has GHC 40,000 of her own money which she used to purchase gas cylinder and stove, refrigerator, microwave, and blender and so forth. If she had invested the money in a treasury bill she would earn GHC 4,000 per year. She has been offered GHC 15,000 per year to work as a cook in AIT. She said her entrepreneurial abilities are worth GHC 3,000 per year. Total annual revenue from pottery sales is $72,000. Calculate Dhorkas’ explicit and implicit costs and


3 the find her accounting and economic profits. Explain why the implicit cost will constitute Dhorkas’ normal profit


1
Expert's answer
2021-05-27T10:49:45-0400

Explicit cost= 12000+5000 +20000

=GHC 37000

Implicit cost= 15000+3000+4000

=GHC 22000

Accounting profit= Total revenue - explicit cost

= 72000 - 37000

= 35000

Economic profit = total revenue - total cost

=72000- 37000- 22000

= 13000

The implicit cost will constitute normal profit because the normal profit does not Include the actual spending of money.